The 1-year $51 options on Nest stock are priced at $1.49 for the call and $1.27 for the put. The annual risk-free rate is 3 percent. What is the per share price of the underlying stock? (Hint: use the put-call parity to determine the share price)

$51.00

$50.25

$50.67

$48.86

$49.73
Answer #1

According to the Put-Call Parity;

C + [X / (1 + r)] = P + S

$1.49 + [$51 / 1.03] = $1.27 + S

S = $1.49 + $49.51 - $1.27 = $49.73

Hence, Option "E" is correct.